OpenStudy (anonymous):

The taxi fare in a city is $3 and then $0.40 for every kilometre travelled. A taxi fare is $9. How far has the taxi travelled? (I have the answer but dont know how to work it out)

OpenStudy (anonymous):

ok first you need to site up the Q mathematically 3+0.4*D where D is the distance travelled [ km]

OpenStudy (anonymous):

D=x, since x more common 3+0.4x=9 solving for x 0.4x=9-3 x=6/.4=15 km

OpenStudy (anonymous):

I hope that makes it clear for you.

OpenStudy (anonymous):

why do we equate 3+0.4x = 9?

OpenStudy (anonymous):

okay.... the fare is 3$ up front..... it has nothing to do with distance

OpenStudy (anonymous):

and 0.4 is the rate per km ...

OpenStudy (anonymous):

so adding these together should give the total charge

OpenStudy (anonymous):

ohh okay